# DICE: Staleness-Centric Optimizations for Parallel Diffusion MoE Inference

### v2

@@ 5.2 Main Results
Speedups and memory analysis. DICE effectively achieves a 1.2$\times$ acceleration through expert parallelism while maintaining the quality. As demonstrated in Figure 9, DICE consistently exhibits significant speedups over expert parallelism across various batch sizes and image resolutions on both DiT-MoE-XL and G, reaching a maximum of 1.26$\times$ at batch size 32. Indeed, interweaved parallelism alone achieves better image quality compared to DistriFusion while maintaining similar latency, and DICE further trades latency for even higher image quality in a cost-effective manner.
@@
Class-Conditional ImageNet 256$\times$256
@@
Displaced Expert Parallelism 27.61 15.42 111.05 0.55 1.28$\times$
DICE 15.13 14.52 166.32 0.66 1.20$\times$
@@
Class-Conditional ImageNet 256$\times$256
@@
Displaced Expert Parallelism 15.27 13.44 156.82 0.64 1.33$\times$
DICE 8.60 12.3 205.05 0.71 1.24$\times$
@@ 5.3 Experiments on Fewer Steps
The 50-step setting was selected to align with the publicly available DiT-MoE codebase, which hardcodes this parameter. To ensure broader applicability, we performed additional comparative experiments using 10 and 20 sampling steps.
Quality results. From Tables 2 and 3, it is clear that even in fewer steps, DICE continues to exhibit superior performance. In terms of quality, DICE achieves FID scores of 15.13 and 8.60, which are significantly lower than those of DistriFusion and the displaced expert parallelism. These results further demonstrate that our method achieves acceleration while incurring minimal quality loss.
Speedups and memory analysis. DICE continues to gain favorable acceleration at reduced steps. As evidenced in Tables 2 and 3, it maintains a speedup exceeding 1.2×1.2$\times$1.2 ×. In comparison to DistriFusion, DICE consistently exhibits a notable advantage in memory efficiency, whereas DistriFusion encounters out-of-memory (OOM) in these setups.
Class-Conditional ImageNet 256$\times$256
@@ 5.4 Ablation Study
Selective Synchronization. Our analysis indicated that deeper layers are more sensitive to staleness. Ablation results (Tab. 4) confirm that synchronizing only the deeper half (Deep) yields the best image quality, effectively trading a small latency increase for substantial gains.
@@ 6 Conclusion
We propose DICE, a staleness-centric optimization framework that accelerates MoE-based diffusion model inference through three core innovations: Interweaved Parallelism for reducing step-level staleness, Selective Synchronization for selectively synchronizing staleness-sensitive layers, and Conditional Communication, a token-aware strategy that dynamically adjusts communication frequency. Together, these techniques reduce memory usage, enhance inference efficiency, and achieve up to 1.26×1.26$\times$1.26 × speedup over expert parallelism with minimal quality loss. We anticipate that DICE will pave the way for scalability and efficiency improvements in diffusion model serving.
@@ 7 More Preliminaries
Figure 11: Illustration of an MoE model[5] with one shared experts and two non-shared experts.
@@
Displaced parallelism [20] overlaps communication and computation by asynchronously transmitting activations computed in the current step for use in the next step (Algorithm 2); this prevents blocking (Algorithm 1). As shown in Figure 12, each device sends activations without waiting, continuing computation with slightly outdated activations from the previous step. This approach prevents communication-induced blocking but introduces staleness, as computations rely on slightly outdated activations rather than fresh data.
1 for each layer do
@@ 2       $x\leftarrow\text{attn}(x)$
       $[s,\,idx]\leftarrow\text{router}(x)$ // Compute router scores and assign tokens to experts
@@
3       $[s,\,idx]\leftarrow\text{router}(x)$
@@ 8 Alternative Staggered Batch Solution
Many works split batches to better overlap communication and computation [9, 18]. We initially explored a staggered batch solution, in which each device divides its local batch into multiple sub-batches and processes them in a staggered manner. As shown in Figure 13, this approach can reduce staleness by handling multiple sub-batches.
@@
  • Increased Memory Requirement. Unlike interweaved parallelism—which requires a persistent buffer only for combine and a temporary one for dispatch—the staggered batch solution must maintain persistent buffers for both dispatch and combine, effectively doubling memory usage and increasing the risk of out-of-memory (OOM) errors.
@@
Figure 14: Batch size scaling performance of DiT-MoE on 8 NVIDIA RTX 3080 GPUs.
Figure 15: Image size scaling performance of DiT-MoE on 8 NVIDIA RTX 3080 GPUs.
@@ 2       $x\leftarrow\text{attn}(x)$
3       $[s,\,idx]\leftarrow\text{router}(x)$
@@
[…]
@@
1 for token-expert pair $(i,e)$ at step $t$ do
@@ 10 Discussion
Applicability to NVLink. While our experiments are conducted on PCIe-connected GPUs, DICE remains applicable when MoE models are served under NVLink and InfiniBand-based multi-node deployments [39]. In such settings, all-to-all communication can contribute up to 76% of total inference latency [39], suggesting that DICE might offer even greater benefits in these environments.
Integration with Existing Expert Parallelism Optimizations. Existing expert parallelism optimizations—such as expert shadowing [9], topology-aware routing [9, 19], and affinity-based methods [39]—address orthogonal challenges; respectively: load balancing, network topology, and expert placement. These techniques are potentially integrable with DICE rather than alternatives, and combining them could further enhance overall efficiency.
@@ 11 Additional Experiments
All-to-All Blocking Latency in Expert Parallelism. We measure all-to-all latency on the same hardware as Sec. 5. Table 5 reveals a significant bottleneck inherent in synchronous Expert Parallelism for DiT-MoE models: the all-to-all communication overhead. The experimental data indicates that, across all tested configurations, all-to-all communication time significantly surpasses half of the total inference time. Furthermore, a pronounced upward trend is observed in this percentage as the batch size increases, reaching a strikingly high 79.2%. This underscores the necessity of our proposed approach to mitigate this communication bottleneck and achieve substantial inference acceleration.
@@
Speedups and memory analysis. The extra experiments were conducted on 8 NVIDIA RTX 3080 (20GB) GPUs [^1] Conducted on GPUs rented from the AutoDL (https://www.autodl.com/home), which may potentially include non-standard or modified configurations. connected via PCIe, powered by 96 vCPU Intel(R) Xeon(R) Platinum 8352V CPU @ 2.10GHz, with 384GB of memory.
The results shown in Figure 14 and 15. Similar to the results on the NVIDIA RTX 4090 GPUs, DICE also demonstrates significant speedup across various batch sizes and image resolutions, achieving a maximum speedup of 23% at batch size 32. The speedup on the 3080 GPUs is slightly lower than on the 4090 GPUs under the same configuration (Which is 26.1%). This may be due to the 3080’s lower computational performance, which reduces the impact of communication overhead on overall inference time.

### v4

@@ 5.2 Main Results
[…]
@@
Class-Conditional ImageNet $\times256$
@@
[…]
@@
Class-Conditional ImageNet $\times256$
@@
[…]
@@ 5.3 Experiments on Fewer Steps
To ensure broader applicability, we performed additional comparative experiments using 10 and 20 sampling steps.
Quality results. Table 2 and 3 show that even in fewer steps, DICE continues to exhibit better performance. In terms of quality, DICE achieves FID scores of 15.13 and 8.60, which are lower than those of DistriFusion and the displaced expert parallelism. These results further demonstrate that our method achieves acceleration while incurring minimal quality loss.
Speedups and memory analysis. DICE continues to gain favorable acceleration at reduced steps. As evidenced in Tables 2 and 3, it maintains a speedup exceeding $1.2\times$. In comparison to DistriFusion, DICE consistently exhibits a notable advantage in memory efficiency, whereas DistriFusion encounters out-of-memory (OOM) in these setups.
Class-Conditional ImageNet $\times256$
@@ 5.4 Ablation Study
Selective Synchronization. Our analysis indicated that deeper layers are more sensitive to staleness. Ablation results (Table 4) confirm that synchronizing only the deeper half (Deep) yields the best image quality, effectively trading a small latency increase for substantial gains.
@@ 6 Conclusion
We propose DICE, a staleness-centric optimization framework that accelerates MoE-based diffusion model inference through three core innovations: Interweaved Parallelism for reducing step-level staleness, Selective Synchronization for selectively synchronizing staleness-sensitive layers, and Conditional Communication, a token-aware strategy that dynamically adjusts communication frequency. Together, these techniques reduce memory usage, enhance inference efficiency, and achieve up to $1.26\times$ speedup over expert parallelism with minimal quality loss. We anticipate that DICE will pave the way for scalability and efficiency improvements in diffusion model serving.
Acknowledgment
This work is supported in part by the National Key Research and Development Project of China (Grant No. 2023YFF0905502), National Natural Science Foundation of China (Grant No. 92467204 and 62472249), and Shenzhen Science and Technology Program (Grant No. JCYJ20220818101014030 and KJZD20240903102300001).
[…]
@@ 7 More Preliminaries
Figure 11: Illustration of an MoE model[ditmoe] with one shared experts and two non-shared experts.
@@
Displaced parallelism [distri] overlaps communication and computation by asynchronously transmitting activations computed in the current step for use in the next step (Algorithm 2). this prevents blocking (Algorithm 1). As shown in Figure 12, each device sends activations without waiting, continuing computation with slightly outdated activations from the previous step. This approach prevents communication-induced blocking but introduces staleness, as computations rely on slightly outdated activations rather than fresh data.
1for each layer do
@@ 2  $x\leftarrow\text{attn}(x)$
  $s,\,idx]\leftarrow\text{router}(x)$ // Compute router scores and assign tokens to experts
@@
3  $s,\,idx]\leftarrow\text{router}(x)$
@@ 8 Alternative Staggered Batch Solution
Many works split batches to better overlap communication and computation [fastermoe, lina]. We initially explored a staggered batch solution, in which each device divides its local batch into multiple sub-batches and processes them in a staggered manner. As shown in Figure 13, this approach can reduce staleness by handling multiple sub-batches.
@@
  • Increased Memory Requirement. Unlike interweaved parallelism—which requires a persistent buffer only for combine and a temporary one for dispatch—the staggered batch solution must maintain persistent buffers for both dispatch and combine, effectively doubling memory usage and increasing the risk of out-of-memory.
@@ 2  $x\leftarrow\text{attn}(x)$
3  $s,\,idx]\leftarrow\text{router}(x)$
@@
  $\text{prev\_layer\_x}_{d}\leftarrow\text{wait}(\text{prev\_layer\_h}_{d})$ // Wait for previous layer’s dispatch result from current step
  $\text{prev\_layer\_x}_{e}\leftarrow\text{prev\_layer\_expert}(\text{prev\_layer\_x}_{d})$ // Process tokens using previous layer’s experts
  $\text{prev\_layer\_h}_{c}\leftarrow\text{async\_combine}(\text{prev\_layer\_x}_{e})$ // Launch non-blocking combine for previous layer’s outputs
@@
1for token-expert pair $(i,e)$ at step $t$ do
@@ 10 Discussion
Applicability to NVLink. While our experiments are conducted on PCIe-connected GPUs, DICE remains applicable when MoE models are served under NVLink and InfiniBand-based multi-node deployments [interlayer]. In such settings, all-to-all communication can contribute up to 76% of total inference latency [interlayer], suggesting that DICE might offer even greater benefits in these environments.
Integration with Existing Expert Parallelism Optimizations. Existing expert parallelism optimizations—such as expert shadowing [fastermoe], topology-aware routing [fastermoe, locmoe], and affinity-based methods [interlayer]—address orthogonal challenges; respectively: load balancing, network topology, and expert placement. These techniques are potentially integrable with DICE rather than alternatives, and combining them could further enhance overall efficiency.
@@ 11 Additional Experiments
All-to-All Blocking Latency in Expert Parallelism. We measure all-to-all latency on the same hardware as Section 5. Table 5 reveals a significant bottleneck inherent in synchronous Expert Parallelism for DiT-MoE models: the all-to-all communication overhead. The experimental data indicates that, across all tested configurations, all-to-all communication time significantly surpasses half of the total inference time. Furthermore, a pronounced upward trend is observed in this percentage as the batch size increases, reaching a strikingly high 79.2%. This underscores the necessity of our proposed approach to mitigate this communication bottleneck and achieve substantial inference acceleration.